A 290-mL gas sample contains argon and nitrogen at a temperature of 66 ∘C. The total pressure of the sample is 738 mmHg and the partial pressure of argon is 262 mmHg.

What mass of nitrogen is present in the sample?
Answer #1

262 = mole fraction of Ar x 738

mole fraction of Ar = 0.355

mole fraction of N2 = 0.645

volume = 0.290 L

T = 66 + 273 = 339 K

Ptotal = 738 / 760 = 0.971 atm

P V = n R T

0.971 x 0.290 = n x 0.0821 x 339

n = 0.0101

total moles = 0.0101

moles of N2 / total moles = mole fraction of N2

moles of N2 / 0.0101 = 0.645

moles of N2 = 6.5 x 10^-3

mass of N2 = 6.5 x 10^-3 x 28 = 0.182 g

mass of N2 = 0.182 g